OpenStudy (anonymous):

Tanya drives to work every day and passes two independently operated traffic lights. The probability that both lights are red is 0.55. The probability that the first light is red is 0.69. What is the probability that the second light is red, given that the first light is red?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

I think that the requested probability \(p\), has to check the subsequent condition: \(0.69 \cdot p=0.55\)

OpenStudy (anonymous):

Then would I divide?

OpenStudy (michele_laino):

yes! That equation expresses the fact that the two events are independent

OpenStudy (anonymous):

um 0.797?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

I need 1 more?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

A survey of 1,200 men and women asked, "Do you earn over $75,000 per year?" The table below shows the responses for males and females: Male Female Total Income over $75,000 585 485 1,070 Income below $75,000 65 65 130 Total 650 550 1200 Based on these data, are "being female" and "earning over $75,000" independent events?

OpenStudy (anonymous):

Based on these data, are "being female" and "earning over $75,000" independent events? No, P(being female | the person earns over $75,000) = P(being female) No, P(being female | the person earns over $75,000) ≠ P(being female) Yes, P(being female | the person earns over $75,000) = P(being female) Yes, P(being female | the person earns over $75,000) ≠ P(being female)

OpenStudy (michele_laino):

I'm sorry, I don't know this answer

OpenStudy (anonymous):

Pam is playing a board game and rolls two number cubes. Let A = {the sum of the number cubes is even}, and let B = {the sum of the number cubes is divisible by 3}. List the outcomes in A ∪ B. {2, 3, 4, 6, 8, 9, 10, 12}

OpenStudy (anonymous):

Thats the answer i got:)

OpenStudy (michele_laino):

please wait, I'm working on your question...

OpenStudy (anonymous):

kk:)

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Papa's Italian Restaurant has collected data about customer sauce orders. It calculated that P(marinara) = 0.74, P(pesto) = 0.62, and P(marinara or pesto) = 0.90. Determine the P(marinara and pesto).

OpenStudy (anonymous):

Is it 0.46?

OpenStudy (michele_laino):

we have: \(0.74+0.62=0.90+x\) from which \(x=0.46\) so, your answer is correct!

OpenStudy (anonymous):

Thanks so much for all your help!